OpenStudy (anonymous):

claudia has a cake recipe that requires 4 cups of flour, 4 eggs, 1 cup of sugar, 2 cups of milk and 1/2 cup of cocoa powder . claudia only has 3 eggs. she decides to alter the recipe to use only 3 eggs. how much of each ingredient will she need?

OpenStudy (texaschic101):

this is all about ratios.. ratio of eggs to flour = 4/4 or just 1....so for each egg, there is 1 cup of flour. So using 3 eggs, you would use 3 cups of flour. ratio of eggs to sugar is 4/1 4/1 = 3/x --- (4 eggs to 1 cup sugar = 3 eggs to x cups of sugar cross multiply 4x = 3 x = 3/4 so if 3 eggs are used, then 3/4 cups of sugar is used. ratio of eggs to milk is 4/2 4/2 = 3/x --(4 eggs to 2 cups milk = 3 eggs to x cuos of milk) cross multiply 4x = 6 x = 6/4 x = 1 1/2 so if 3 eggs are used, then 1 1/2 cups milk is used. ratio of eggs to cocoa powder is 4/(1/2) 4/(1/2) = 3/x -- (4 eggs to 1/2 cup cocoa = 3 eggs to x cups cocoa) cross multiply 4x = 3/2 x = (3/2) / 4 x = 3/2 * 1/4 x = 3/8 so if 3 eggs were used, then 3/8 cup of cocoa is used. To sun it up, the ingredients when used with 3 eggs are as follows : 3 eggs 3 cups of flour 3/4 cups of sugar 1 1/2 (or 1.5) cups of milk 3/8 cups of cocoa powder